OpenStudy (anonymous):

Find the domain of f(x,y)=sin^-1[(x-y)^2]. Do this mathematically and then describe the domain in words or with a sketch.

OpenStudy (anonymous):

(x-y)^2<1

OpenStudy (anonymous):

can you explain any?

OpenStudy (anonymous):

you cannot take the arcsine of a number less than -1 or greater than one so your answer is \[-1\leq (x-y)^2\leq 1\] but of course \[(x-y)^2\geq 0\]

OpenStudy (anonymous):

this says the distance between x and y has to be less than or equal to one, and the graph is the region between two parallel lines here is a picture http://www.wolframalpha.com/input/?i=%28x-y%29^2%3C1